VEGAS Team Meeting

Dial in Number: 877-937-8961

Passcode: 9445091#

Attendees: Ray, Randy, Adam, Dan, Amanda, Sam, Mark, Paul, Marty

Agenda:

Whole group:

  1. Status of BOFs for Modes 2 and 3 (H16k)

GB only

  1. "top hat" problem
  2. Review of "known problems" from: https://safe.nrao.edu/wiki/bin/view/CICADA/VegasStatusUpdate
  3. Differed Mode 10-19 tests until Joe returns
  4. IF target levels
  5. Plans for test time

Meeting notes:

Whole group:

  1. Status of BOFs for Modes 2 and 3 (H16k)
    • Dan to check with Hong and send an update on anticipated delivery
    • Complete ripple tests when new Mode 2/3 BOFs are available/tested

GB only

  1. "top hat" problem
    • Awaiting Joe
  2. Review of "known problems" from: https://safe.nrao.edu/wiki/bin/view/CICADA/VegasStatusUpdate
    • high Tsys on first integration problem waiting for Joe
    • Some reports of ~20% varying Tsys across channels - team suspects calibration
  3. Differed Mode 10-19 tests until Joe returns
    • Agreed
  4. IF target levels
    • The table for the IF Rack balancing will be different since the attenuator settings in the Converter Rack will now have different values for VEGAS than for the Spectrometer
    • Anish / Galen - to confirm how the IF levels into VEGAS were set in the first place?
  5. New issues
    1. "Bank E appeared to be a `sick' bank. It either dropped data during acquisition or failed to stop acquisition at the end of the scan in a few cases."
      • Manager has capability to capture this item Anish reported
    2. Calibration (ADC)
      • Calibrations no longer fails - should we change manager to remove use of configuration file for values?
      • Paul to do an analysis of the issue and report back to group with MR
  6. Plans for test time
    • Unscheduled request(s): none until new BOFs are available
    • Tuesday
    • Wednesday
    • Thursday
    • Monday (8/18)
      • None at this time

Other

  1. Please make sure to properly code your timesheets for the current time period and going forward.
    • Software engineers should be using a GBT ops software CM&E account - Account 893252
    • Electronic engineers should be using a GBT ops electronics CM&E account - Account 432510
    • Commissioning activities should be getting charged to hardware support and testing - Account 432510
    • Scientists assisting with shared-risk commissioning should be charging an observer support account
      • 432540 for the most part. Accounts 142020 (GB-based staff) and 142010 (CV-based staff) for when you are helping with overarching observing support (not problem solving, VEGAS specific)

-- MartyBloss - 2014-08-11
Topic revision: r1 - 2014-08-11, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback